Lab Report Calculations for The spectrum of Hydrogen - Physics i Belum mm The Spectrum of Hydrogen Purpose: Determine the key beng constant wavelenghts of H-spectrum lines in the Buliner s
Answer #1

Using the given formula for λ(mm)=(d sinθ )/m =sinθ/(2*295.2)

and λ(nm) = 106 * λ(mm)

And comparing this wavelength with λ(balmer)=3646.14*(n2/(n2-4)) Angstrom to get the value of n, the table fills out as:

1/ện*n) 1/ (cm) 6 0.027778 24359.86 8 Y 01 02 A3 Oi=0-00 sin(Oi) A(nm) MA) Nearest Abalmer n 144 144.05 144.05 144.0333 14.03

Now, the plot of 1/n2 vs 1/λ(cm)

1/n2 vs 1/1(cm) 30000 25000 20000 1/(cm) 15000 y = -110573x + 27428 10000 5000 0 0 0.02 0.04 0.06 0.08 0.1 0.12 1/n?

The slope comes out to be -110573.

Now, the negative of slope gives the Rydberg Constant.

so R=110573 cm-1

Actual value of Rydberg constant =109737cm-1

%error = {|(110573-109737)/109737|}*100=0.76%